# A Superconducting Single-Atom Phonon Laser

### v1

@@ CONCLUSION
This article demonstrates the experimental realization of a phononic analog to the optical single-atom laser. Our experiment consists of a superconducting single-atom, realized using a transmon style qubit coupled resonantly to an HBAR mode. When driving on resonance with the HBAR mode, the intrinsic non-linearity of the qubit-phonon coupling generates a highly excited phonon state — the phonon laser. A key feature of this experiment is the population inversion of a two-level atom rather than the parametric instability driving phonon excitations. Moreover, the phonon mode is confined in a bulk longitudinal mode within a sapphire substrate. The bulk nature may allow the phonon laser’s integration with additional on-chip architectures, such as color centers or quantum dots.
Our results have demonstrated the successful generation of a coherent phonon laser, and we have further demonstrated a counterintuitive feature of phonon lasers, an upper instead of a lower lasing threshold Ashhab et al. (2009). The size of the coherent lasing state achieved in this work was limited by both the qubit and phonon linewidth. However, the primary limiting factor was the phonon linewidth. Decreasing the decay rates will reduce the upper lasing threshold power and increase the peak lasing amplitude. Future studies could include a linear probe of the phonon mode to probe the phonon statistics. Such an on-chip phonon laser promises to provide a highly coherent source of phonons, which have applications from sensing to quantum information processing.
@@ ACKNOWLEDGMENTS
Authors contributions C.A.P. performed experiments, theoretical modelling, conceptualization, and wrote the manuscript. W.J.M.F. fabricated the device and performed experiments. V.A.S.V.B. performed theoretical modelling. A.M. provided supervision and funding acquisition. G.A.S. provided supervision, conceptualization and funding acquisition. Competing interests: The authors declare no competing interests.
@@
Supplementary Information
I Materials and Methods
I.1 Device Fabrication
@@
Once fabrication on both chips was done, the HBAR chip was diced into 8x2 mm chips. The HBAR chip was then flipped on top of the qubit chip with the AlN layer facing down. Using probe needles, the AlN disks were aligned with the transmon antennas. Once aligned, the probe needles held down the chips in position while a tapered fiber was used to apply two-component epoxy (Loctite EA 3430) on the sides of the top chip; see Fig. S1. After the epoxy was cured, the chip was wire-bonded and installed onto the baseplate of the dilution refrigerator.
I.2 Measurement Setup
@@
Figure S2: Schematic of the measurement setup. To the left is the general Dilution refrigerator wiring setup. Outside the refrigerator, we used two different setups. In (b), we show our Two-tone spectroscopy configuration used for qubit spectroscopy. Here, we sent a weak continuous wave tone (readout resonator probe) from the vector network analyzer (VNA) port one and a second continuous wave tone (qubit drive) from the VNA port 3. These two signals are combined by using a directional coupler before entering the dilution refrigerator. The signal from the dilution refrigerator goes through a room-temperature amplifier before it goes to port 2 of the VNA. In (c), we show are ”time domain” setup. Here, we replaced output port 3 of the VNA with a signal generator (SG) to provide the drive tone. A switch (Taylor MW switch) together with a pulse generator (PG, Rigol DG1022) is placed between the directional coupler and the signal generator.
All measurements were performed within a dilution refrigerator operating at a base temperature $T\sim 20$ mK. A schematic of the dilution refrigerator setup and the room-temperature electronics are shown in Fig. S2. The device was mounted on the mixing chamber plate of the dilution refrigerator and connected to a set of coaxial cables. The device was measured in transmission, with the resonators coupled in a ’notch’-style geometry. The output signals went into a cryogenic HEMT (High Electron Mobility Transistor) amplifier (LNF-LNC4-8A), followed by additional room-temperature amplification (Miteq AFS3-04000800-07-10P-4). The input line was attenuated at each stage to reduce the electron temperature and the thermal radiation at the input port of our device. A total of 48 dB of attenuation was used, plus any additional attenuation from the coaxial cables.
@@
For the gated two-tone measurement, port one of the VNA was set up just as in the two-tone measurement. An external signal generator generated the qubit drive tone. The qubit drive tone was passed through an RF switch before being combined with the readout tone using a directional coupler. The RF switch was triggered using a pulse generator at a 3 kHz repetition rate. The signal generator was set to a 22.5% duty cycle, so the qubit drive was off for 75 $\mu$s per trace. The VNA was synchronously triggered by the signal generator allowing for 65536 trace averages to be performed (maximum setting). Each trace averaged measurement was repeated 75 additional times to improve the signal-to-noise ratio further.
@@ II.1 System Hamiltonian and Master Equation
The Lamb-shifted qubit frequency is $\tilde{\Omega}_{\rm q}=\omega_{\rm q}+\chi$. Given the system’s parameters, we will discard the qubit-mediated beam-splitter term $\frac{g_{\rm qb}}{g_{\rm qc}}\chi(\hat{a}^{\dagger}\hat{b}+\hat{a}\hat{b}^{\dagger})\hat{\sigma}_{\rm z}$. Furthermore, in the two-tone spectroscopic setup, we can retain only the probe term for the cavity and only the drive term for the qubit. With such approximations, we have
@@
where $\Delta_{\rm b}=\omega_{\rm d}-\omega_{\rm b}$, $\Delta_{\rm q}=\omega_{\rm d}-\tilde{\Omega}_{\rm q}$, and $\Delta_{\rm r}=\omega_{\rm p}-\omega_{\rm r}+\chi$. We have defined the readout cavity detuning $\Delta_{\rm r}$ to included the Lamb shift, however, this decision is arbitrary. The density matrix of the system $\rho$ has dynamics described by the master equation
@@
II.2 Qubit Two-Tone Spectroscopy
@@
II.3 Master Equation Simulations
@@
where $\delta$ calibrates the room-temperature power to the corresponding value of $\mathcal{P}_{\rm d}$ accounting for all losses and multiplicative factors. The value of $\delta$ was determined by matching the simulation to the power-broadened qubit spectrum at multiple qubit drive powers. A set of qubit spectra is shown in Fig. S4 where it can be seen that the master equation simulation is in excellent agreement capturing the qubit asymmetry at low power, and the power-broadened qubit spectrum at higher drive powers. It should be noted the measured transmission signal $S_{21}$ is proportional to the qubit population $\langle\hat{\sigma}_{\rm z}\rangle$ only if the appropriate signal quadrature is measured. Here this was determined by calculating the rotation angle that minimized the signal in the out-of-phase quadrature. To confirm, we compared our single-trace two-tone data with a direct measurement of the readout resonator frequency at several qubit drive powers. At the highest drive powers, we observed a slight deviation resulting from signal mixing into the out-of-phase quadrature. However, this was at the drive powers above the self-quenching threshold and resulted in a slight mismatch between our experimental and simulated two-tone traces.
@@
Following the calibration of the qubit spectrum, the phonon mode was included in the master equation simulation. The value of the coupling rate $g_{\rm qb}$ and the phonon linewidth $\gamma_{\rm b}$ were determined by performing a fit to the time domain gated two-tone measurements. The fit was determined by fitting both the ringdown (as shown in the main text) and also the ring-up of the qubit. The best-fit values were then compared at multiple drive powers; see Fig. S5(a). We extracted a value of the qubit phonon coupling of $g_{\rm qb}/2\pi=162$ kHz and a phonon linewidth of $\gamma_{\rm b}/2\pi=6.81$ kHz. From this set of simulations, the steady-state phonon population and phonon statistics can be estimated for multiple drive powers. We observe a good agreement between the numerical simulation and the ring-up data and an excellent agreement between the ringdown data for all powers. Deviations in the ring-up simulations likely result from higher-order nonlinearities we are not considering in our model.

### v2

@@ CONCLUSION
This article demonstrates the experimental realization of a phononic analog to the optical single-atom laser. Our experiment consists of a superconducting single-atom, realized using a transmon style qubit coupled resonantly to an HBAR mode. When driving on resonance with the HBAR mode, the intrinsic non-linearity of the qubit-phonon coupling generates a highly excited phonon state — the phonon laser. A key feature of this experiment is the driven two-level atom rather than the parametric instability driving phonon excitations. Moreover, the phonon mode is confined in a bulk longitudinal mode within a sapphire substrate. The bulk nature may allow the phonon laser’s integration with additional on-chip architectures, such as color centers or quantum dots.
Our results have demonstrated the successful generation of a coherent phonon laser, and we have further demonstrated a counterintuitive feature of phonon lasers, an upper instead of a lower lasing threshold Ashhab et al. (2009). The size of the coherent lasing state achieved in this work was limited by both the qubit and phonon linewidth. However, the primary limiting factor was the phonon linewidth. Decreasing the decay rates will reduce the upper lasing threshold power and increase the peak lasing amplitude. Moreover, it should be possible to construct a single-atom microwave photon laser using a superconducting qubit weakly coupled to a microwave resonator in conventional cQED. Future studies could include a linear probe or use a higher-order transition of the transmon to drive and measure the phonon statistics. Such a cQAD-compatible phonon laser promises to provide a highly coherent source of phonons, which have applications including sensing to quantum information processing and the generation of high-displacement, Schrödinger cat states.
@@ ACKNOWLEDGMENTS
Authors contributions C.A.P. performed experiments, theoretical modelling, conceptualization, and wrote the manuscript with input from all authors. W.J.M.F. fabricated the device and performed experiments. V.A.S.V.B. performed theoretical modelling. A.M. provided supervision and funding acquisition. G.A.S. provided supervision, conceptualization and funding acquisition.
Ethics declarations:
Competing interests: The authors declare no competing interests.
@@
Methods
Device Fabrication
@@ Methods
Once fabrication on both chips was done, the HBAR chip was diced into 8x2 mm chips. The HBAR chip was then flipped on top of the qubit chip with the AlN layer facing down. Using probe needles, the AlN disks were aligned with the transmon antennas. Once aligned, the probe needles held down the chips in position while a tapered fiber was used to apply two-component epoxy (Loctite EA 3430) on the sides of the top chip; see the supplementary information. After the epoxy was cured, the chip was wire-bonded and installed onto the baseplate of the dilution refrigerator.
Measurement Setup
@@
All measurements were performed within a dilution refrigerator operating at a base temperature $T\sim 20$ mK. A schematic of the dilution refrigerator setup and the room-temperature electronics are shown in the supplementary information. The device was mounted on the mixing chamber plate of the dilution refrigerator and connected to a set of coaxial cables. The device was measured in transmission, with the resonators coupled in a ’notch’-style geometry. The output signals went into a cryogenic HEMT (High Electron Mobility Transistor) amplifier (LNF-LNC4-8A), followed by additional room-temperature amplification (Miteq AFS3-04000800-07-10P-4). The input line was attenuated at each stage to reduce the electron temperature and the thermal radiation at the input port of our device. A total of 48 dB of attenuation was used, plus any additional attenuation from the coaxial cables.
@@
For the gated two-tone measurement, port one of the VNA was set up just as in the two-tone measurement. An external signal generator generated the qubit drive tone. The qubit drive tone was passed through an RF switch before being combined with the readout tone using a directional coupler. The RF switch was triggered using a pulse generator at a 3 kHz repetition rate. The signal generator was set to a 22.5% duty cycle, so the qubit drive was off for 75 $\mu$s per trace. The VNA was synchronously triggered by the signal generator, allowing for 65536 trace averages to be performed (maximum setting). Each trace averaged measurement was repeated 75 additional times to improve the signal-to-noise ratio further.
Supplementary Information
I Device
Figure S2: Schematic of the measurement setup. To the left is the general Dilution refrigerator wiring setup. Outside the refrigerator, we used two different setups. In (b), we show our Two-tone spectroscopy configuration used for qubit spectroscopy. Here, we sent a weak continuous wave tone (readout resonator probe) from the vector network analyzer (VNA) port one and a second continuous wave tone (qubit drive) from the VNA port 3. These two signals are combined using a directional coupler before entering the dilution refrigerator. The signal from the dilution refrigerator goes through a room-temperature amplifier before it goes to port 2 of the VNA. In (c), we show our ”time domain” setup. Here, we replaced output port 3 of the VNA with a signal generator (SG) to provide the drive tone. A switch (Taylor MW switch) together with a pulse generator (PG, Rigol DG1022) is placed between the directional coupler and the signal generator.
@@ II.1 System Hamiltonian and Master Equation
The Lamb-shifted qubit frequency is $\tilde{\Omega}_{\rm q}=\omega_{\rm q}+\chi$. Given the system’s parameters, we will discard the qubit-mediated beam-splitter term $\frac{g_{\rm qb}}{g_{\rm qc}}\chi(\hat{a}^{\dagger}\hat{b}+\hat{a}\hat{b}^{\dagger})\hat{\sigma}_{\rm z}$, as it is several orders of magnitude smaller than $g_{\rm qb}$. Furthermore, in the two-tone spectroscopic setup, we can retain only the probe term for the cavity and only the drive term for the qubit. With such approximations, we have
@@
where $\Delta_{\rm b}=\omega_{\rm d}-\omega_{\rm b}$, $\Delta_{\rm q}=\omega_{\rm d}-\tilde{\Omega}_{\rm q}$, and $\Delta_{\rm r}=\omega_{\rm p}-\omega_{\rm r}+\chi$. We have defined the readout cavity detuning $\Delta_{\rm r}$ to include the Lamb shift; however, this decision is arbitrary. The density matrix of the system $\rho$ has dynamics described by the master equation
@@
II.2 Mean field theory for the steady-state phonon population
We can eliminate the microwave cavity mode in (S5) following the procedure outlined in Gambetta et al. (2008). The procedure corresponds to a displacement of the cavity mode conditioned on the qubit state, followed by a partial trace of the cavity mode under the assumption of no occupation of the cavity fluctuations. More details will be given in a future paper. The procedure yields the following master
$\partial_{t}\varrho=-\frac{i}{\hbar}[\hat{\mathcal{H}}_{\rm eff},\varrho]+\gamma_{\rm{b}}\mathcal{L}[\hat{b}]\varrho+\Gamma_{1}\mathcal{L}[\hat{\sigma}_{-}]\varrho+\frac{\tilde{\Gamma}_{\phi}}{2}\mathcal{L}[\hat{\sigma}_{\rm{z}}]\varrho,$ (S6)
where $\varrho$ is the phonon-qubit density matrix, and $\tilde{\Gamma}_{\phi}$ is the qubit dephasing including the read-out cavity induced dephasing Gambetta et al. (2008). The effective Hamiltonian reads
^$\frac{\hat{\mathcal{H}}_{{\rm{eff}}}}{\hbar}=-\Delta_{\rm{b}}\hat{b}^{\dagger}\hat{b}-\frac{\tilde{\Delta}_{\rm{q}}}{2}\hat{\sigma}_{\rm{z}}+g_{\rm{qb}}\left(\hat{b}\hat{\sigma}_{+}+\hat{b}^{\dagger}\hat{\sigma}_{-}\right)+\varepsilon_{\rm{d}}\left(\hat{\sigma}_{+}+\hat{\sigma}_{-}\right).$ (S7)
The qubit detuning $\tilde{\Delta}_{q}$ also considers the modification due to the read-out cavity.
The total qubit dephasing and the qubit detuning are given by
$\tilde{\Gamma}_{\phi}$ $=\Gamma_{\phi}+\Gamma_{\phi,\rm{cav}}(t),$ (S8)
$\tilde{\Delta}_{\rm{q}}$ $=\Delta_{\rm{q}}-\omega_{\rm{q,cav}}(t),$
$\omega_{\rm{q,cav}}(t)(t)$ $=2\chi{\rm{Re}}[\alpha_{\rm{g}}(t)\alpha_{e}^{*}(t)],$
$\Gamma_{\phi,\rm{cav}}(t)$ $=2\chi{\rm{Im}}[\alpha_{\rm{g}}(t)\alpha_{\rm{e}}^{*}(t)].$
Here, $\alpha_{{}_{\rm{e(g)}}}(t)$ are the cavity amplitude if the qubit is in the excited (ground) state, given by
$\partial_{t}\alpha_{{\rm{e}}}$ $=\left[i(\Delta_{{\rm{r}}}-\chi)-\frac{\kappa}{2}\right]\alpha_{{\rm{e}}}-i\epsilon_{{\rm{p}}},$ (S9)
$\partial_{t}\alpha_{{\rm{g}}}$ $=\left[i(\Delta_{{\rm{c}}}+\chi)-\frac{\kappa}{2}\right]\alpha_{{\rm{g}}}-i\epsilon_{{\rm{p}}}.$
Since the cavity decay $\kappa$ is significantly larger than the other decays in the system, we consider for now on that the cavity is always in its steady state, such that the amplitudes $\alpha_{\rm{e(g)}}$ are given by
$\alpha_{\rm{e}}$ $=\frac{i\epsilon_{\rm{p}}}{i(\Delta_{{\rm{r}}}-\chi)-\frac{\kappa}{2}},$ (S10)
$\alpha_{\rm{g}}$ $=\frac{i\epsilon_{\rm{p}}}{i(\Delta_{{\rm{r}}}+\chi)-\frac{\kappa}{2}}.$
From the master equation (S6), we can then obtain the following equations for $\langle\hat{b}\rangle=b,$ $\langle\hat{\sigma}_{-}\rangle=s_{-}$, $\langle\hat{\sigma}_{\rm{z}}\rangle=s_{\rm{z}}$:
$\partial_{t}b$ $=\left(i\Delta_{\rm{b}}-\frac{\gamma_{\rm{b}}}{2}\right)b(t)-ig_{\rm{qb}}s_{-}(t),$ (S11)
$\partial_{t}s_{-}$ $=\left(i\tilde{\Delta}_{\rm{q}}-\tilde{\gamma}_{2}\right)s_{-}(t)+ig_{\rm{qb}}b(t)s_{\rm{z}}(t)+i\varepsilon_{\rm{q}}s_{\rm z}(t),$
$\partial_{t}s_{\rm{z}}$ $=2is_{-}(t)\left(g_{\rm{qb}}b^{*}(t)+\varepsilon_{\rm{d}}\right)-2is_{-}^{*}(t)\left(g_{\rm{qb}}b(t)+\varepsilon_{\rm{d}}\right)-\gamma_{1}(s_{\rm{z}}(t)+1),$
which assume the mean field approximations $\langle\hat{b}\hat{\sigma}_{\rm{z}}\rangle\approx\langle\hat{b}\rangle\langle\hat{\sigma}_{\rm{z}}\rangle=bs_{\rm{z}}$ and $\langle\hat{b}^{\dagger}\hat{\sigma}_{-}\rangle\approx\langle\hat{b}^{\dagger}\rangle\langle\hat{\sigma}_{-}\rangle=b^{*}s_{-}$. The steady-state phonon number is given by $\bar{n}_{b}=|\bar{b}|^{2}$, where $\bar{b}$ is the steady-state of the phonon amplitude $b(t)$. Such steady-state is given by
$\bar{n}_{\rm{b}}=|\bar{b}|^{2}=\frac{g_{\rm{qb}}^{2}\varepsilon_{\rm{d}}^{2}\bar{s}_{\rm{z}}^{2}}{\left(\Delta_{\rm{b}}\tilde{\gamma}_{2}+\tilde{\Delta}_{\rm{q}}\frac{\gamma_{\rm{b}}}{2}\right)^{2}+\left(\frac{\gamma_{\rm{b}}\tilde{\gamma}_{2}}{2}-\Delta_{\rm{b}}\tilde{\Delta}_{\rm{q}}-g_{\rm{qb}}^{2}\bar{s}_{\rm{z}}\right)^{2}},$ (S12)
where the steady-state $\bar{s}_{\rm{z}}$ reads
$\bar{s}_{\rm z}=-\frac{\gamma_{1}}{\gamma_{1}+4\frac{\tilde{\gamma}_{2}}{\tilde{\Delta}_{\rm{q}}^{2}+{\tilde{\gamma}_{2}}^{2}}|g_{\rm{qb}}\bar{b}+\varepsilon_{\rm{d}}|^{2}}.$ (S13)
The equation for the phonon number (S12) is a non-linear equation for $\bar{b}$ that has to be solved numerically. The results for parameters in correspondence with the experiment are shown in Fig.  S3, in which we can see a good agreement with the full master equation simulation.
Figure S3: Steady-state mean-field phonon number as a function of the qubit drive power. The red line depicts the result obtained with the mean-field solution of Eq. (S14), and the gray points depict the results from simulations of the full master equation (S5). The parameters used are given in Table 1 in correspondence with the experimental parameters of the main text.
We can further manipulate equation (S14) as
$\bar{n}_{\rm{b}}=|\bar{b}|^{2}=\frac{g_{\rm{qb}}^{2}\varepsilon_{\rm{d}}^{2}}{\left(\Delta_{\rm{b}}\frac{\gamma_{\rm{eff}}}{2}+\Delta_{\rm{eff}}\frac{\gamma_{\rm{b}}}{2}\right)^{2}+\left(\frac{\gamma_{\rm{b}}\gamma_{\rm{eff}}}{4}-\Delta_{\rm{b}}\Delta_{\rm{eff}}+g_{\rm{qb}}^{2}\right)^{2}},$ (S14)
where we have defined the effective decay $\gamma_{\rm{eff}}$ and detuning $\Delta_{\rm{eff}}$ as
$\frac{\gamma_{\rm{eff}}}{2}$ $\equiv\frac{\tilde{\gamma}_{2}}{-\bar{s}_{\rm{z}}}=\frac{\tilde{\gamma}_{2}}{\gamma_{1}}\left(\gamma_{1}+4\frac{\tilde{\gamma}_{2}}{\tilde{\Delta}_{\rm{q}}^{2}+{\tilde{\gamma}_{2}}^{2}}|g_{\rm{qb}}\bar{b}+\varepsilon_{\rm{d}}|^{2}\right),$ (S15)
$\Delta_{\rm{eff}}$ $\equiv\frac{\tilde{\Delta}_{\rm{q}}}{-\bar{s}_{\rm{z}}}=\frac{\tilde{\Delta}_{\rm{q}}}{\gamma_{1}}\left(\gamma_{1}+4\frac{\tilde{\gamma}_{2}}{\tilde{\Delta}_{\rm{q}}^{2}+{\tilde{\gamma}_{2}}^{2}}|g_{\rm{qb}}\bar{b}+\varepsilon_{\rm{d}}|^{2}\right).$
From such a mean-field perspective, the steady-state of the phonon mode behaves similarly as if it would be coupled to a coherently driven cavity with the nonlinear decays and detunings given by Eq.  (S15). We notice that as the power is increased, the effective decay $\gamma_{\rm{eff}}$ increases, a consequence of the qubit power broadening. Such an effect yields the eventual decoupling between the qubit and the phonon, similar to an inverse Purcell effect. Nevertheless, such a power broadening is not the only physical mechanism that plays a role in the build-up of the phonon steady-state. In fact, both the effective decay and detuning depend on the phonon amplitude $\bar{b}$, a non-linearity that can not be discarded for the experimental parameters. To show this effect, we plot in Fig. S4 the mean-field solution for $n_{\rm{b}}$ (red curve) and the corresponding curve by setting $g_{\rm{qb}}=0$ in Eqs. (S15) (dashed black curve). The latter still captures the inverse Purcell effect due to the qubit power broadening but yields half of the maximum phonon occupancy, indicating the importance of the phonon nonlinearity stemming from the phonon-qubit coupling and the intrinsic two-level nature of the qubit.
Figure S4: Steady-state mean-field phonon number as a function of the qubit drive power. The red line depicts the result obtained with the mean-field solution of Eq. (S14), while the black dashed line depicts the result obtaining by discarding the phonon nonlinearity in eqs. (S15) The parameters used are given in Table 1 in correspondence with the experimental parameters of the main text.
II.3 Qubit Two-Tone Spectroscopy
@@
II.4 Master Equation Simulations
@@
where $\delta$ calibrates the room-temperature power to the corresponding value of $\mathcal{P}_{\rm d}$ accounting for all losses and multiplicative factors. The value of $\delta$ was determined by matching the simulation to the power-broadened qubit spectrum at multiple qubit drive powers. A set of qubit spectra is shown in Fig. S6 where it can be seen that the master equation simulation is in excellent agreement capturing the qubit asymmetry at low power, and the power-broadened qubit spectrum at higher drive powers. It should be noted the measured transmission signal $S_{21}$ is proportional to the qubit population $\langle\hat{\sigma}_{\rm z}\rangle$ only if the appropriate signal quadrature is measured. Here, this was determined by calculating the rotation angle that minimized the signal in the out-of-phase quadrature. To confirm, we compared our single-trace two-tone data with a direct measurement of the readout resonator frequency at several qubit drive powers. At the highest drive powers, we observed a slight deviation resulting from signal mixing into the out-of-phase quadrature. However, this was at the drive powers above the self-quenching threshold and resulted in a slight mismatch between our experimental and simulated two-tone traces.
@@
Following the calibration of the qubit spectrum, the phonon mode was included in the master equation simulation. The value of the coupling rate $g_{\rm qb}$ and the phonon linewidth $\gamma_{\rm b}$ were determined by performing a fit to the time domain gated two-tone measurements. The fit was determined by fitting both the ringdown (as shown in the main text) and also the ring-up of the qubit. Here, the device was allowed to thermalize before the gated-ringdown measurement, weakly probing the readout resonator continually. At the time $t=0$, the qubit drive is switched on for $175\mu$s, allowing the phonon moded to ring-up. At $t=175\mu$s, the qubit drive is switched off, and the phonon ringdown is observed, as described in the main text. The best-fit values were compared at multiple drive powers; see Fig. S7(a). We extracted a value of the qubit phonon coupling of $g_{\rm qb}/2\pi=162$ kHz and a phonon linewidth of $\gamma_{\rm b}/2\pi=6.81$ kHz. From this set of simulations, the steady-state phonon population and phonon statistics can be estimated for multiple drive powers. We observe a good agreement between the numerical simulation and the ring-up data and an excellent agreement between the ringdown data for all powers. Deviations in the ring-up simulations likely result from higher-order nonlinearities we are not considering in our model.
@@
III Simplified Single-Atom Laser Model
III.1 Considerations of Qubit Anharmonicity
We modeled the single-atom laser using a two-level approximation; however, under strong drives, it is important to consider the higher levels of the transmon qubit. We can instead model the qubit as a Kerr oscillator with a large anharmonicity $\alpha$. Therefore, we can re-write the Hamiltonian in the form
^$\frac{\hat{\mathcal{H}}_{\rm mod}}{\hbar}=-\Delta_{\rm q}\hat{c}^{\dagger}\hat{c}+\frac{\alpha}{2}\hat{c}^{\dagger}\hat{c}^{\dagger}\hat{c}\hat{c}-\Delta_{\rm b}\hat{b}^{\dagger}\hat{b}+g_{\rm qb}(\hat{b}^{\dagger}\hat{c}+\hat{b}\hat{c}^{\dagger})+\varepsilon_{\rm d}(\hat{c}^{\dagger}+\hat{c}).$ (S21)
Where we have replaced the two-level raising and lowering operators with bosonic raising and lowering operators $\hat{c}^{(\dagger)}$. If we calculate the number of steady-state phonons comparing the Kerr oscillator model to the two-level system model, we find at powers below and near the upper lasing threshold, the population is slightly larger in the Kerr oscillator model; see Fig. S8. There exists a discrepancy at high drive powers; however, this was not observed within the experiment since we likely could not apply a large enough drive without heating the sample. Thus, we are confident that modeling our experiment as a two-level system is accurate, especially for calculating the upper lasing threshold. Moreover, this was required for numerical efficiency. The two-level system simulation in Fig. S8 ran for approximately 20 minutes, whereas considering only three levels of the Kerr oscillator increased the simulation time to approximately 10 hours.
Figure S8: Two-Level verses Anharmonic Oscillator Simulations. (A) The simulated phonon population when considering the transmon qubit a two-level system versus considering three levels of the Kerr oscillator.
For numerical efficiency, we have constructed a simplified model where we have removed the coupling to the readout resonator. Studying Eq. S3, the readout resonator broadens the qubit and causes a constant frequency offset due to steady-state photons in the cavity. In this section, we aim to explore the properties of the single-atom laser. The Hamiltonian of our simplified model is given by
^$\frac{\hat{\mathcal{H}}_{\rm mod}}{\hbar}=-\frac{\Delta_{\rm q}}{2}\hat{\sigma}_{\rm z}-\Delta_{\rm b}\hat{b}^{\dagger}\hat{b}+g_{\rm qb}(\hat{b}^{\dagger}\hat{\sigma}_{-}+\hat{b}\hat{\sigma}_{+})+\varepsilon_{\rm d}(\hat{\sigma}_{+}+\hat{\sigma}_{-}),$ (S22)
and we choose to have the qubit and phonon resonant for the simulation. Moreover, we have increased intrinsic qubit linewidth $\Gamma_{\rm q}=1.5$ MHz to account for the lack of readout resonator broadening. We also increased the phonon linewidth $\gamma_{\rm b}=25$ kHz to reduce the size of the Hilbert space for numerical efficiency.
III.2 Phonon Phase Coherence
First, we aim to investigate if the drive tone seeds the phase of the phonon mode. In the model described in Ref. Ashhab et al. (2009) the phase of the lasing mode is random and results in a distribution in the IQ plane. However, if we examine the phase space distribution of the phonon mode in our model, we find that the phase of the drive tone instead sets the phase; see Fig. S9. This results from the weak hybridization between the qubit and phonon mode. When the drive is switched on, the weak hybridization enables the seeding of the phase of the phonon mode prior to the onset of lasing. This is followed by the stimulated emission of phonons, which produces a large coherent state with a set phase. The ability to set the phase of the phonon mode is important for applications such as the generation of Schrödinger cat states.
Figure S9: Phonon Lasing Phase Seeding. (A) Phonon population as a function of qubit drive power. Colored points correspond to the Wigner plots (B-E). (B-E) Wigner distribution plots corresponding to the points in (A). The phase of the lasing state is set by the drive phase; here, the drive phase was $\pi/4$. The plots are shown starting from the lowest power (B) to the largest power (E).
III.3 Phonon Mode Anharmonicity
Finally, the coupling between the qubit and phonon mode raises the question of how much anharmonicity is inherited by the phonon mode. In the strong coupling limit, the phonon mode will inherit half of the qubit anharmonicity and, thus, would be unable to lase since its Hilbert space is truncated to a single excitation. From the transparency window measurement, we observe little frequency shift of the phonon mode, suggesting the mode is well approximated as a harmonic oscillator and not a Duffing oscillator. However, again, we can utilize our simplified model to explore the inherited anharmonicity by directly driving the phonon mode in the simulation, bypassing all lasing dynamics. We find that at a drive strength that generates, on average, fifty phonons, similar to what was observed in the experiment, the phonon mode has a negligible frequency shift. Suggesting that in the weak hybridization limit, the phonon mode can be treated to a high degree as a linear harmonic oscillator.